The manager of the Petroco Service Station wants to forecast the demand for unleaded gasoline next month so that the proper number of gallons can be ordered from the distributor. The owner has accumulated the following data on demand for unleaded gasoline from sales during the past 10 months: MAT540 Homework Week 4 Page 2 of 5 Month Gasoline Demanded (gal.) October 775 November 835 December 605 January 450 February 600 March 700 April 820 May 925 June July 1500 1200 a. Compute an exponential smoothed forecast, using an α value of 0.4 b. Compute the MAD.
Answer #1

The exponential smoothed forecast is obtained using the following equation

F = F:-1 + a(At-1-F:-1)

Forecast for month of October = 775

Forecast for November = 775 + 0.4 ( 775 - 775) = 775

Forecast for December = 775 + 0.4 ( 835 - 775 ) = 799

Forecast for January = 799 + 0.4 ( 605 - 799) = 721.4

Forecast for February = 721.4 + 0.4 ( 450 - 721.4) = 612.84

Forecast for March = 612.84 + 0.4 ( 600 - 612.84 ) = 607.70

Forecast for April = 607.704 + 0.4 ( 700 - 607.704 ) = 644.62

Forecast for May = 644.62 + 0.4 ( 820 - 644.62 ) = 714.77

Forecast for June = 714.77 + 0.4 ( 925 - 714.77 ) = 798.86

Forecast for July =  798.86 + 0.4 ( 1500 - 798.86 ) = 1079.32

Month Gasoline demanded Forecasted demand
October 775 775
November 835 775
December 605 799
January 450 721.4
February 600 612.84
March 700 607.7
April 820 644.62
May 925 714.77
June 1500 798.86
July 1200 1079.32

---------------------------------------------------------------------------------------

To calculate the mean absolute deviation ( MAD)

60 271.4 Gasoline demanded Forecasted demand Deviation = Actual - forecast Absolute deviatic 775 775 835 775 60 605 799 -194

Mean absolute deviation = 1837.97 \div 10

Mean absolute deviation = 183.80